# Constructing an Optimal Behavior Basis for the Option Keyboard

### v1

@@ Appendix A Proofs
Appendix B Beyond Linear Rewards
@@
We start by arguing that, in many cases, the linear reward assumption does not pose any limitations. As discussed by Barreto et al., (2020), when both $\mathcal{S}$ and $\mathcal{A}$ are finite, we can recover any possible reward function by defining $d=|\mathcal{S}|^{2}\times|\mathcal{A}|$ features $\phi_{i}$, each being an indicator function associated with a specific transition $(s,a,s^{\prime})$. This result shows that it is possible to define reward features $\bm{\mathbf{\phi}}\in\mathbb{R}^{|\mathcal{S}|^{2}\times|\mathcal{A}|}$, such that every task in $\mathcal{M}$ is linearly expressible. A challenge, however, is to find alternative features with this property, but with dimension $d\ll|\mathcal{S}|$ Touati and Ollivier, (2021).
@@
Unfortunately, the OK may be unable to solve all tasks in the family $\mathcal{M}_{\text{sd-linear}}^{\bm{\mathbf{\phi}}}$, even when having access to a set of policies $\Pi_{k}$ forming a CCS.
Proposition B.1.
There exists a task $M\in\mathcal{M}_{\text{sd-linear}}^{\bm{\mathbf{\phi}}}$ such that no meta-policy $\omega:\mathcal{S}\to\mathcal{Z}$ and set of policies $\Pi$ results in $\pi^{\text{OK}}_{\omega}(s;\Pi)$ being optimal with respect to $M$.
@@
Assume a family of MDPs $M\in\mathcal{M}_{\text{sd-linear}}^{\bm{\mathbf{\phi}}}$ with states, actions, transitions, and features defined as in Figure 6.
@@
Let a task $M\in\mathcal{M}_{\text{sd-linear}}^{\bm{\mathbf{\phi}}}$ be defined via the following reward function: $r(s_{1})=\bm{\mathbf{\phi}}(s_{1})\cdot[1,0]=0,r(s_{2})=\bm{\mathbf{\phi}}(s_{2})\cdot[-1,-1]=-3,r(s_{3})=\bm{\mathbf{\phi}}(s_{3})\cdot[-1,-1]=-3,r(s_{4})=\bm{\mathbf{\phi}}(s_{4})\cdot[1,1]=2$, with optimal policy $\pi^{*}(s_{0})=a_{4}$ whose SF vector is $\bm{\mathbf{\psi}}^{\pi^{*}}=[1,1]$. We now show that there is no meta-policy $\omega(s)$ that results in $\pi^{\text{GPI}}(s;\omega(s))$ being equal to the optimal policy $\pi^{*}$. This is proved by noticing the impossibility of the following equality:
@@
Appendix C Training the OK Meta-Policy
In Alg. 3, we show the pseudocode for the algorithm that optimizes the OKB meta-policy $\omega$ for a given set of tasks $\mathcal{W}^{\text{sup}}$.
@@ Appendix C Training the OK Meta-Policy
1:  Input: Meta-policy $\omega$, weight support $\mathcal{W}^{\text{sup}}$, base policies $\Pi$.
2:  Initialize replay buffer $\mathcal{B}$
3:  Let $\bm{\mathbf{\psi}}^{\omega}(s,\bm{\mathbf{z}},\bm{\mathbf{w}})$ be the critic of the meta-policy $\omega$
4:  $\bm{\mathbf{w}}_{0}\sim\mathcal{W}^{\text{sup}}$; $S_{0}\sim\mu$
5:  for $t=0,...,T$ do
6:     if $S_{t}$ is terminal then
7:        $\bm{\mathbf{w}}_{t}\sim\mathcal{W}^{\text{sup}}$
8:        $S_{t}\sim\mu$
9:     end if
10:     $\bm{\mathbf{z}}_{t}\leftarrow\omega(S_{t},\bm{\mathbf{w}})$
11:     $\bm{\mathbf{z}}_{t}\leftarrow\bm{\mathbf{z}}_{t}+\mathrm{clip}(\epsilon,-0.5,0.5)$, where $\epsilon\sim\mathcal{N}(0,0.2^{2})$  $\triangleright$ Exploration clipped Gaussian noise, as in TD3
12:     $\bm{\mathbf{z}}_{t}\leftarrow\bm{\mathbf{z}}_{t}/||\bm{\mathbf{z}}_{t}||_{2}$
13:     $A_{t}\leftarrow\pi^{\text{GPI}}(S_{t},\bm{\mathbf{z}}_{t};\Pi)$  $\triangleright$ Follow OK policy
14:     Execute $A_{t}$, observe $S_{t+1}$, and $\bm{\mathbf{\phi}}_{t}$
15:     Add $(S_{t},\bm{\mathbf{z}}_{t},\bm{\mathbf{\phi}}_{t},S_{t+1})$ to $\mathcal{B}$
16:     Update $\bm{\mathbf{\psi}}^{\omega}$ by minimizing $\mathbb{E}_{(s,\bm{\mathbf{z}},\bm{\mathbf{\phi}},s^{\prime})\sim\mathcal{B},\bm{\mathbf{w}}\sim\mathcal{W}^{\text{sup}}}\left[\left(\bm{\mathbf{\psi}}^{\omega}(s,\bm{\mathbf{z}},\bm{\mathbf{w}})-(\bm{\mathbf{\phi}}+\gamma\bm{\mathbf{\psi}}^{\omega}(s^{\prime},\omega(s^{\prime},\bm{\mathbf{w}}),\bm{\mathbf{w}}))\right)^{2}\right]$
17:     Update $\omega$ via the policy gradient $\nabla_{\bm{\mathbf{z}}}\bm{\mathbf{\psi}}^{\omega}(s,\bm{\mathbf{z}},\bm{\mathbf{w}})\cdot\bm{\mathbf{w}}|_{\bm{\mathbf{z}}=\omega(s,\bm{\mathbf{w}})}\nabla_{\omega}\omega(s,\bm{\mathbf{w}})$
18:  end for
19:  Return $\omega$
Appendix D Checking the Condition for OK Optimality
Our practical implementation of the OKB test the condition in Thm. 4.2 by randomly sampling from a replay buffer, =$\mathcal{B}{=}\{(s_{i},a_{i},\bm{\mathbf{\phi}}(s_{i},a_{i}),s^{\prime}_{i},)\}_{i=1}^{n}$, which contains experiences observed during the training of the base policies in $\Pi_{k}$. Given a candidate task $\bm{\mathbf{w}}\in\mathcal{C}$ (line 12 of Alg. 1), we compute the mean positive advantage of $\omega$ as:
@@ Appendix D Checking the Condition for OK Optimality
where $\hat{A}_{\bm{\mathbf{w}}}^{\omega}(s_{i},a_{i})=\bm{\mathbf{\phi}}(s_{i},a_{i})\cdot\bm{\mathbf{w}}+\gamma\bm{\mathbf{\psi}}^{\omega}(s^{\prime}_{i},\omega(s^{\prime}_{i}))\cdot\bm{\mathbf{w}}-\bm{\mathbf{\psi}}^{\omega}(s_{i},\omega(s_{i}))\cdot\bm{\mathbf{w}}$. In line 21 of Alg. 1, we select the task $\bm{\mathbf{w}}\in\mathcal{C}$ with the highest value for Eq. (21). Intuitively, we select the task $\bm{\mathbf{w}}$ for which the OK can improve the most its performance. By adding $\pi_{\bm{\mathbf{w}}}$ to $\Pi_{k}$ and retraining $\omega$, the OK will be more expressive in the subsequent iteration. We highlight that OKB theoretical guarantees (e.g., Thm. 4.3) are independent of the heuristic used to select the task $\bm{\mathbf{w}}\in\mathcal{C}$ (line 21), and other strategies could be used instead.
Appendix E Experiments Details
@@
The USFA $\bm{\mathbf{\psi}}(s,a,\bm{\mathbf{w}})$ used for encoding the base policies $\Pi_{k}$ were modeled with multi-layer perceptron (MLP) neural networks with 4 layers with 256 neurons. We use an ensemble of $10$ neural networks, similar to Chen et al., (2021), and compute the minimum value over two randomly sampled elements when computing the Bellman update targets. We used Leaky ReLU activation functions and layer normalization for improved training stability. We used the same neural network architectures and hyperparameters for OKB and the competing baselines for fairness of comparison.
@@
The meta-policy $\omega(s,\bm{\mathbf{w}})$ was modeled with an MLP with 3 layers with 256 neurons. We employed the techniques introduced by Bhatt et al., (2024), i.e., batch normalization and removal of target networks, which increased the training efficiency. We used Adam (Kingma and Ba, , 2015) as the first-order optimizer used to train all neural networks with mini-batches of size $256$.
When training the base policies, we used $\varepsilon$-greedy exploration with a linearly decaying schedule. For training the meta-policy, we added a clipped Gaussian noise (see line 11 of Alg. 3), as done in other actor-critic algorithms, e.g., TD3 (Fujimoto et al., , 2018).
@@
To generate sets of test tasks $\mathcal{W}^{\prime}\subset\mathcal{W}$ given different values of $d$, we employed the method introduced by Takagi et al., (2020) available on pymoo (Blank and Deb, , 2020), which produces uniformly-spaced weight vectors in the simplex, $\mathcal{W}$.
E.1 Corner Weights.
Below, we define the concept of corner weights used by OKB (Alg. 1) and OK-LS (Alg. 2), as defined in previous works (Roijers, , 2016; Alegre et al., , 2022).
Definition E.1.
@@
where $\bm{\mathbf{V}}^{+}$ is a matrix whose rows store the elements of $\Psi$ and is augmented by a column vector of $-1$’s. Each vector $\bm{\mathbf{x}}{=}(w_{1},...,w_{d},v_{\bm{\mathbf{w}}})$ in $P$ is composed of a weight vector and its scalarized value.
To compute corner weights, as in Def. E.1, we used pycddlib (https://github.com/mcmtroffaes/pycddlib/) implementation of the Double Description Method (Motzkin et al., , 1953) to efficiently enumerate the vertices of the polyhedron $P$.
E.2 Domains

### v2

@@ Appendix A Proofs
A.5 Conditions for OKB’s Behavior Basis to be Strictly Smaller Than the CCS
Transfer learning techniques dating back to the early 1990s often leverage the principles of compositionality and modularity—the idea that related tasks can be solved by reusing parts of previously learned solutions. [^8] These principles underpin a long line of research on hierarchical reinforcement learning, temporal abstraction, macro-actions, skill discovery, and studies in neuroscience and cognitive science exploring how animal behavior is organized hierarchically. Our formal analysis of when OKB’s behavior basis is strictly smaller than the CCS draws on the principles of compositionality and modularity—the idea that related tasks can be solved by combining reusable components (“sub-behaviors”) from previously acquired policies. We show that identifying these conditions reduces to analyzing the cardinality of a specific Set Cover problem.
Intuitively, the OKB only requires a behavior basis as large as the CCS in degenerate cases where the solution to all possible tasks involves executing at least one type of behavior that does not occur in the solutions to any other tasks. Such cases are fundamentally at odds with the assumptions that make transfer learning viable. In contrast, as long as policies in the CCS (which solve related tasks drawn from a shared distribution) produce similar behaviors in shared regions of the state space, the OKB matches the expressiveness of the CCS while requiring strictly fewer policies.
Consider a transfer learning scenario where agents must solve tasks drawn from some distribution, and each task’s solution can be expressed as a (possibly stochastic, closed-loop) sequence of sub-behaviors—often referred to as options, skills, policy modules, controllers, motor primitives, or sub-policies. For example, task $\tau_{1}$ might require sub-behaviors $o_{1}$, $o_{2}$, $o_{5}$, and $o_{6}$ (opening doors, driving, turning off lights, and finding exits), while task $\tau_{2}$ might require a different combination of sub-behaviors—some of which it may share with $\tau_{1}$, since both come from the same distribution.
Assume a distribution of related tasks that induce a CCS whose policies use some subset of the available sub-behaviors, $O=\{o_{1},\dots,o_{N}\}$, with $\pi_{i}$ drawing on $O_{i}\subseteq O$. The question is: What is the smallest subset of CCS policies that together cover all sub-behaviors in $O$? Formally, we seek the smallest collection $\{\pi_{j},\pi_{k},\pi_{r},\dots\}$ such that $O_{j}\cup O_{k}\cup O_{r}\cup\dots=O$. This reduces to the classic Set Cover problem and corresponds to identifying the minimum number of CCS policies containing all sub-behaviors that might be required to solve any tasks from the distribution of interest.
Let $\mathrm{SC}\subseteq\mathrm{CCS}$ denote the solution to this Set Cover problem. By Proposition 4.4, if the solution to a novel task consists of switching between policies in a behavior basis, then a corresponding optimal meta-policy exists that sequences such policies. To formally characterize when the OKB is as expressive as a CCS while relying on a strictly smaller behavior basis, we need to understand when $|\mathrm{SC}|<|\mathrm{CCS}|$.
In general, a Set Cover solution will have $|\mathrm{SC}|=|\mathrm{CCS}|$ only in degenerate cases where no CCS policy can be discarded without leaving some sub-behavior uncovered. One such case arises when every policy in the CCS contains at least one unique sub-behavior—a behavior that never contributes to solving any other possible task. This situation is arguably at odds with the assumptions of transfer learning. Only in such cases—for example, when every possible task uniquely requires a specialized sub-behavior that no other task ever uses—would $|\mathrm{SC}|=|\mathrm{CCS}|$.
As long as some sub-behaviors are shared between elements of the CCS, the set of policies needed to cover all sub-behaviors (i.e., $|\mathrm{SC}|$) will be strictly smaller than the CCS. In such common transfer learning settings, the OKB is as expressive as a CCS while requiring strictly fewer policies.
Appendix B Transforming General Linear Reward MDPs into Convex Reward MDPs
Let $\mathcal{M}_{\text{lin}}^{\bm{\mathbf{\phi}}}\subseteq\mathcal{M}$ be the (possibly infinite) set of MDPs associated with all linearly expressible reward functions. This set, commonly studied in the SFs literature, can be defined as
$\mathcal{M}_{\text{lin}}^{\bm{\mathbf{\phi}}}\triangleq\{(\mathcal{S},\mathcal{A},p,r_{\bm{\mathbf{w}}},\mu,\gamma)\mid r_{\bm{\mathbf{w}}}=\bm{\mathbf{\phi}}\cdot\bm{\mathbf{w}}\},$
where $\bm{\mathbf{w}}\in\mathbb{R}^{d}$ is a $d$-dimensional weight vector that may include negative elements.
Standard methods for computing convex coverage sets (CCS) require reward functions to be expressed as convex combinations of features, i.e., using weights in the $d$-dimensional simplex, $\Delta^{d}$:
$\Delta^{d}=\left\{\bm{\mathbf{w}}\in\mathbb{R}^{d}\mid w_{i}\geq 0,\sum_{i=1}^{d}w_{i}=1\right\}.$
To use methods that learn a CCS—typically designed for convex reward functions—to solve all linear tasks in $\mathcal{M}_{\text{lin}}^{\bm{\mathbf{\phi}}}$, we first need to show that any MDP in $\mathcal{M}_{\text{lin}}^{\bm{\mathbf{\phi}}}$ can be transformed into an equivalent one where rewards are expressed as convex combinations over a new set of features. In other words, given a family of MDPs $\mathcal{M}_{\text{lin}}^{\bm{\mathbf{\phi}}}$, we show how to construct a corresponding family of MDPs with convex rewards (defined over a transformed feature space $\tilde{\bm{\mathbf{\phi}}}$) that induces the same set of optimal policies. This family of MDPs can be defined as follows:
$\mathcal{M}_{conv}^{\tilde{\bm{\mathbf{\phi}}}}\triangleq\{(\mathcal{S},\mathcal{A},p,r_{\bm{\mathbf{w}}},\mu,\gamma)\mid r_{\bm{\mathbf{w}}}=\tilde{\bm{\mathbf{\phi}}}\cdot\bm{\mathbf{w}},\bm{\mathbf{w}}\in\Delta^{2d}\},\ \text{where}\ \tilde{\bm{\mathbf{\phi}}}(s,a)=\begin{bmatrix}\bm{\mathbf{\phi}}(s,a)\\ -\bm{\mathbf{\phi}}(s,a)\end{bmatrix}\in\mathbb{R}^{2d}.$ (18)
We now show that this transformed family of MDPs induces the same set of optimal policies as $\mathcal{M}_{\text{lin}}^{\bm{\mathbf{\phi}}}$. As a result, learning a CCS for $\mathcal{M}_{\text{conv}}^{\tilde{\bm{\mathbf{\phi}}}}$ allows us to recover the optimal policy for any task in $\mathcal{M}_{\text{lin}}^{\bm{\mathbf{\phi}}}$. Let $r(s,a)=\bm{\mathbf{\phi}}(s,a)\cdot\bm{\mathbf{w}}$ be an arbitrary linear reward function with a weight vector $\bm{\mathbf{w}}\in\mathbb{R}^{d}$. One can decompose $\bm{\mathbf{w}}$ into its non-negative and non-positive components as follows:
$\bm{\mathbf{w}}=\bm{\mathbf{w}}^{+}-\bm{\mathbf{w}}^{-},\quad\text{where}\quad w^{+}_{i}=\max(w_{i},0),\quad w^{-}_{i}=\max(-w_{i},0).$
We can now define an augmented feature vector $\tilde{\bm{\mathbf{\phi}}}$:
$\tilde{\bm{\mathbf{\phi}}}(s,a)=\begin{bmatrix}\bm{\mathbf{\phi}}(s,a)\\ -\bm{\mathbf{\phi}}(s,a)\end{bmatrix}\in\mathbb{R}^{2d},$
and a non-negative weight vector $\tilde{\bm{\mathbf{w}}}$:
$\tilde{\bm{\mathbf{w}}}=\begin{bmatrix}\bm{\mathbf{w}}^{+}\\ \bm{\mathbf{w}}^{-}\end{bmatrix}\in\mathbb{R}_{\geq 0}^{2d}.$
This allows us to rewrite the original reward function as follows:
$r(s,a)=\bm{\mathbf{w}}\cdot\phi(s,a)=\tilde{\bm{\mathbf{w}}}\cdot\tilde{\phi}(s,a).$
Finally, let us normalize the weights $\tilde{\bm{\mathbf{w}}}$ (e.g., by dividing by their norm) to ensure they lie in the simplex $\Delta^{2d}$. Recall that scaling a reward function by a positive constant does not affect its optimal policy, so this transformation is valid. Since we showed that any linear reward function can be expressed as a convex combination of a transformed feature set—without altering the corresponding optimal policy—it follows that a CCS for $\mathcal{M}_{\text{conv}}^{\tilde{\bm{\mathbf{\phi}}}}$ contains the optimal solutions for all tasks in the original family, $\mathcal{M}_{\text{lin}}^{\bm{\mathbf{\phi}}}$.
Appendix C Beyond Linear Rewards
@@
We start by arguing that, in many cases, the linear reward assumption does not pose any limitations. As discussed by Barreto+2020, when both $\mathcal{S}$ and $\mathcal{A}$ are finite, we can recover any possible reward function by defining $d=|\mathcal{S}|^{2}\times|\mathcal{A}|$ features $\phi_{i}$, each being an indicator function associated with a specific transition $(s,a,s^{\prime})$. This result shows that it is possible to define reward features $\bm{\mathbf{\phi}}\in\mathbb{R}^{|\mathcal{S}|^{2}\times|\mathcal{A}|}$, such that every task in $\mathcal{M}$ is linearly expressible. A challenge, however, is to find alternative features with this property, but with dimension $d\ll|\mathcal{S}|$ (Touati&Ollivier2021).
@@
Unfortunately, the OK may be unable to solve all tasks in the family $\mathcal{M}_{\text{sd-linear}}^{\bm{\mathbf{\phi}}}$, even when given access to a set of policies $\Pi_{k}$ forming a CCS.
Proposition C.1.
There exists a task $M\in\mathcal{M}_{\text{sd-linear}}^{\bm{\mathbf{\phi}}}$ such that no meta-policy, $\omega:\mathcal{S}\to\mathcal{Z}$, and set of policies, $\Pi$, results in $\pi^{\text{OK}}_{\omega}(s;\Pi)$ being optimal with respect to $M$.
@@
Assume a family of MDPs, $M\in\mathcal{M}_{\text{sd-linear}}^{\bm{\mathbf{\phi}}}$, with states, actions, transitions, and features defined as in Figure 6.
@@
Let a task $M\in\mathcal{M}_{\text{sd-linear}}^{\bm{\mathbf{\phi}}}$ be defined via the following reward function: $r(s_{1})=\bm{\mathbf{\phi}}(s_{1})\cdot[1,0]=0,r(s_{2})=\bm{\mathbf{\phi}}(s_{2})\cdot[-1,-1]=-3,r(s_{3})=\bm{\mathbf{\phi}}(s_{3})\cdot[-1,-1]=-3,r(s_{4})=\bm{\mathbf{\phi}}(s_{4})\cdot[1,1]=2$, with optimal policy $\pi^{*}(s_{0})=a_{4}$ and whose SF vector is $\bm{\mathbf{\psi}}^{\pi^{*}}=[1,1]$. We now show that there is no meta-policy $\omega(s)$ that results in $\pi^{\text{GPI}}(s;\omega(s))$ being equal to the optimal policy $\pi^{*}$. This is proved by noticing the impossibility of the following equality:
@@
Appendix D Training the OK Meta-Policy
Alg. 3 (Train-OK) is the sub-routine for training a meta-policy omega under an actor-critic training framework, given the selected tasks by Alg. 2. In Alg. 3, we show how our method optimizes the OKB meta-policy $\omega$ for a given set of tasks $\mathcal{W}^{\text{sup}}$.
1
2
3
Input: Meta-policy $\omega$, weight support $\mathcal{W}^{\text{sup}}$, base policies $\Pi$.
4
5Initialize replay buffer $\mathcal{B}$
6 Let $\bm{\mathbf{\psi}}^{\omega}(s,\bm{\mathbf{z}},\bm{\mathbf{w}})$ be the critic of the meta-policy $\omega$
7 $\bm{\mathbf{w}}_{0}\sim\mathcal{W}^{\text{sup}}$; $S_{0}\sim\mu$
8
9for $t=0,1,2,\dots,T$ do
10 if $S_{t}$ is terminal then
11    $\bm{\mathbf{w}}_{t}\sim\mathcal{W}^{\text{sup}}$
12    $S_{t}\sim\mu$
13
14 $\bm{\mathbf{z}}_{t}\leftarrow\omega(S_{t},\bm{\mathbf{w}})$
 $\triangleright$ Exploration clipped Gaussian noise, as in TD3
15 $\bm{\mathbf{z}}_{t}\leftarrow\bm{\mathbf{z}}_{t}+\mathrm{clip}(\epsilon,-0.5,0.5)$, where $\epsilon\sim\mathcal{N}(0,0.2^{2})$
[…]
 $\triangleright$ Follow OK policy
17 $A_{t}\leftarrow\pi^{\text{GPI}}(S_{t},\bm{\mathbf{z}}_{t};\Pi)$
18
19  Execute $A_{t}$, observe $S_{t+1}$, and $\bm{\mathbf{\phi}}_{t}$
20   Add $(S_{t},\bm{\mathbf{z}}_{t},\bm{\mathbf{\phi}}_{t},S_{t+1})$ to $\mathcal{B}$
21
22  Update $\bm{\mathbf{\psi}}^{\omega}$ by minimizing $\mathbb{E}_{(s,\bm{\mathbf{z}},\bm{\mathbf{\phi}},s^{\prime})\sim\mathcal{B},\bm{\mathbf{w}}\sim\mathcal{W}^{\text{sup}}}\left[\left(\bm{\mathbf{\psi}}^{\omega}(s,\bm{\mathbf{z}},\bm{\mathbf{w}})-\left(\bm{\mathbf{\phi}}+\gamma\bm{\mathbf{\psi}}^{\omega}(s^{\prime},\omega(s^{\prime},\bm{\mathbf{w}}),\bm{\mathbf{w}})\right)\right)^{2}\right]$
23
24  Update $\omega$ via the policy gradient $\nabla_{\bm{\mathbf{z}}}\bm{\mathbf{\psi}}^{\omega}(s,\bm{\mathbf{z}},\bm{\mathbf{w}})\cdot\bm{\mathbf{w}}|_{\bm{\mathbf{z}}=\omega(s,\bm{\mathbf{w}})}\nabla_{\omega}\omega(s,\bm{\mathbf{w}})$
25
26return $\omega$
27
@@ 20   Add $(S_{t},\bm{\mathbf{z}}_{t},\bm{\mathbf{\phi}}_{t},S_{t+1})$ to $\mathcal{B}$
Appendix E Checking the Condition for OK Optimality
Our practical implementation of the OKB test the condition in Thm. 4.2 by randomly sampling from a replay buffer, $\mathcal{B}{=}\{(s_{i},a_{i},\bm{\mathbf{\phi}}(s_{i},a_{i}),s^{\prime}_{i},)\}_{i=1}^{n}$, which contains experiences observed during the training of the base policies in $\Pi_{k}$. Given a candidate task $\bm{\mathbf{w}}\in\mathcal{C}$ (line 12 of Alg. 1), we compute the mean positive advantage of $\omega$ as:
@@ Appendix E Checking the Condition for OK Optimality
where $\hat{A}_{\bm{\mathbf{w}}}^{\omega}(s_{i},a_{i})=\bm{\mathbf{\phi}}(s_{i},a_{i})\cdot\bm{\mathbf{w}}+\gamma\bm{\mathbf{\psi}}^{\omega}(s^{\prime}_{i},\omega(s^{\prime}_{i}))\cdot\bm{\mathbf{w}}-\bm{\mathbf{\psi}}^{\omega}(s_{i},\omega(s_{i}))\cdot\bm{\mathbf{w}}$. In line 14 of Alg. 1, we select the task $\bm{\mathbf{w}}\in\mathcal{C}$ with the highest value for Eq. (21). Intuitively, we select the task $\bm{\mathbf{w}}$ for which the OK can improve the most its performance. By adding $\pi_{\bm{\mathbf{w}}}$ to $\Pi_{k}$ and retraining $\omega$, the OK will be more expressive in the subsequent iteration. We highlight that OKB theoretical guarantees (e.g., Thm. 4.3) are independent of the heuristic used to select the task $\bm{\mathbf{w}}\in\mathcal{C}$ (line 14), and other strategies could be used instead.
Appendix F Experimental Details
@@
The USFAs $\bm{\mathbf{\psi}}(s,a,\bm{\mathbf{w}})$ used for encoding the base policies $\Pi_{k}$ were modeled with multi-layer perceptron (MLP) neural networks with 4 layers with 256 neurons. We use an ensemble of $10$ neural networks, similar to Chen+2021, and compute the minimum value over two randomly sampled elements when computing the Bellman update targets. We used Leaky ReLU activation functions and layer normalization for improved training stability. For a fair comparison, we independently optimized all method-specific hyperparameters (e.g., network architectures, OKB settings, and those of competing baselines) via grid search.
@@
The meta-policy $\omega(s,\bm{\mathbf{w}})$ was modeled with an MLP with 3 layers with 256 neurons. We employed the techniques introduced by Bhatt+2024, i.e., batch normalization and removal of target networks, which increased the training efficiency. We used Adam (Kingma&Ba2015) as the first-order optimizer used to train all neural networks with mini-batches of size $256$.
When training the base policies, we used $\varepsilon$-greedy exploration with a linearly decaying schedule. For training the meta-policy, we added a clipped Gaussian noise (see line 11 of Alg. 3), as done in other actor-critic algorithms, e.g., TD3 (Fujimoto+2018).
@@
To generate sets of test tasks $\mathcal{W}^{\prime}\subset\mathcal{W}$ given different values of $d$, we employed the method introduced by Takagi+2020 available on pymoo (blank&deb2020), which produces uniformly-spaced weight vectors in the simplex, $\mathcal{W}$.
All experiments were performed in a cluster with NVIDIA A100-PCIE-40GB GPUs with 32GB of RAM. Each individual run of each method took approximately 2.5 hours (Item Collection), 6 hours (Minecart), 10 hours (FetchPickAndPlace), and 25 hours (Highway).
F.1 Corner Weights.
Below, we define the concept of corner weights used by OKB (Alg. 1) and OK-LS (Alg. 2), as defined in previous works (Roijers2016; Alegre+2022).
Definition F.1.
@@
where $\bm{\mathbf{V}}^{+}$ is a matrix whose rows store the elements of $\Psi$ and is augmented by a column vector of 1-1’s. Each vector $\bm{\mathbf{x}}{=}(w_{1},...,w_{d},v_{\bm{\mathbf{w}}})$ in $P$ is composed of a weight vector and its scalarized value.
To compute corner weights, as in Def. F.1, we used pycddlib (https://github.com/mcmtroffaes/pycddlib) implementation of the Double Description Method (Motzkin+1953) to efficiently enumerate the vertices of the polyhedron $P$.
F.2 Domains